# Federated Graph Condensation with Information Bottleneck Principles

### v3

@@ Abstract
Graph condensation, which reduces the size of a large-scale graph by synthesizing a small-scale condensed graph as its substitution, has benefited various graph learning tasks. However, existing graph condensation methods rely on centralized data storage, which is unfeasible for real-world decentralized data distribution, and overlook data holders’ privacy-preserving requirements. To bridge this gap, we propose and study the novel problem of federated graph condensation (FGC) for graph neural networks (GNNs). Specifically, we first propose a general framework for FGC, in which we decouple the typical gradient matching process for graph condensation into client-side gradient calculation and server-side gradient matching. In this way, the computational burden on the client side is significantly reduced. Nevertheless, our empirical studies show that under the federated setting, the condensed graph will consistently leak data membership privacy, i.e., the condensed graph during the federated training can be utilized to steal the training data under the membership inference attack (MIA). To tackle this issue, we innovatively incorporate information bottleneck principles into the FGC, which only needs to extract partial node features in one local pre-training step and utilize the features during federated training. Extensive experiments on real-world datasets demonstrate that our framework can consistently protect membership privacy during training. Meanwhile, it can achieve comparable and even superior performance against existing centralized graph condensation and federated graph learning methods.
@@ 1 Introduction
Graph data, such as social networks, transportation networks, and molecular interactions, is ubiquitous in the real world. Graph neural networks (GNNs) have been prevalent for modeling these graphs and achieved remarkable success in recent years (Hamilton, Ying, and Leskovec 2017; Kipf and Welling 2017; Velickovic et al. 2018; Xu et al. 2019). However, GNNs have struggled with real-world large-scale graphs due to memory and computational overheads, which are further exacerbated when the models need to be trained multiple times (e.g., neural architecture search (Zhang et al. 2022) and continual learning (Li and Hoiem 2016)). A natural solution from the data-centric view is graph condensation (GC) (Jin et al. 2022b, a; Liu et al. 2022; Zheng et al. 2023). As shown in Figure 1(a), it reduces large-scale graphs into small ones but keeps the utility of original graphs, which facilitates the storage of graph data and computations of downstream tasks. As a typical technique of GC, gradient matching (Jin et al. 2022b, a) forces the gradients of GNN models trained on the original graph and condensed graph to be the same, maintaining a comparable prediction performance but largely reducing the graph size.
@@
Despite significant progress in GC, existing works all hold a basic assumption that the graph data is centrally stored. In reality, the graph data is distributed among different data holders (clients), i.e., each data holder only owns subgraphs. For example, in healthcare systems, data holders (hospitals) only possess a subset of patients (nodes), their information (attributes), and interactions (links). Due to privacy concerns, data holders are unwilling to share their data and many strict regulations such as GDPR (Voigt and Von dem Bussche 2017) also forbid collecting data arbitrarily, which makes centralized GC unfeasible. Federated graph learning (FGL) (Yao et al. 2023; Wan et al. 2022; Zhang et al. 2021b) has emerged in recent years to endow collaboratively training a global model without exposing local data, as is shown in Figure 1(b). Typically, they are dedicated to training a global model by exchanging encrypted information for recovering missing structures (e.g., cross-client edges and neighbors). However, they still suffer from burdensome computations of traditional GNNs for large-scale graph training, especially in the cross-silo scenario (Gao, Yao, and Yang 2022). For example, if the neural architecture search for FGL is conducted, it inevitably performs federated training multiple times, which exacerbates both computations and communications. Therefore, an unexplored area remains while considering GC in the federated setting.
@@
To bridge this gap, we investigate the novel problem of federated graph condensations (FGC), which aims to collaboratively learn a small graph that contains knowledge from different data sources. Subsequently, the condensed graph can be accessed by all data holders for downstream applications, such as node classifications. It’s non-trivial to develop FGC due to two significant challenges. (1) How to preserve the utility of the condensed graph that is learned from distributed subgraphs? Unlike centralized GC the graph data is integral, the entire graph is distributed across data holders in FGC, as is shown in Figure 1(a) and (b). As a result, some critical cross-client edges are missing and these subgraphs also suffer from severe heterogeneity, which inevitably degrades the utility of the condensed graph. In this regard, a unified framework of utility-preserving FGC is desired. (2) How to preserve the local graph’s membership privacy that may be leaked by the condensed graph? Membership privacy is widely used to measure information leakage through deep learning models about their training data (Nasr, Shokri, and Houmansadr 2019). As is shown in Figure 1(b), typical FGL keeps graph data locally and only learns a global model, thus preserving membership privacy. In contrast, a condensed graph needs to be learned and released to the public in FGC, which may contain membership privacy of the local data. Furthermore, due to the accessibility of the condensed graph, traditional model-based defenses for membership privacy are not feasible. It’s expected to design new mechanisms to protect membership privacy during the whole federated training process.
@@
To tackle these challenges, we endeavor to condense graphs in the federated setting and propose a Federated Graph Condensation framework (FedGC). (1) To preserve data utility, we propose a general framework for FGC, the core of which is that the condensed graph is learned by matching the aggregated gradients from clients. Particularly, the gradients are aggregated in a class-aware weighted manner to tackle data heterogeneity. Furthermore, an encrypted one-step communication between the server and clients is adopted to recover the cross-client neighbors. During the evaluation phase, the model trained on the condensed graph will be fine-tuned by local data for better personalization. (2) To preserve membership privacy, we further propose a novel local graph transformation module with information bottleneck principles (Alemi et al. 2017). We conduct empirical studies and find that the condensed graph will consistently leak membership privacy. To address this, the local graph will be first transformed before condensation by fixing the local graph structure unchanged and extracting sufficient but minimal node features. Then the transformed graph will act as a proxy of the original graph to conduct FGC, thereby protecting the local membership privacy. Besides, to alleviate the performance decay in graph transformation caused by scarce labeled nodes and heterogeneous feature space, we propose a one-time self-training strategy to label nodes and utilize a fixed pre-trained shared model to align the feature space. Finally, we give theoretical and empirical analysis to prove the utility- and privacy-preserving properties of FedGC. The major contributions of this paper are summarized as follows:
@@
  • We conduct extensive experiments on five real-world datasets and show that FedGC outperforms centralized GC and FGL methods, especially in large-scale datasets. Furthermore, FedGC can consistently protect membership privacy during the whole federated training process.
@@ 2 Preliminary
MIA on condensed graphs. Following (Dong, Zhao, and Lyu 2022), we assume an honest-but-curious server is the strong attacker. Since the condensed graph $G^{\prime}$ is publicly accessed, the target models can be arbitrarily selected and trained by attacks, which is different from traditional MIA on graphs (the attacker can only query target models (Olatunji, Nejdl, and Khosla 2021)), as is shown in Figure 2. Therefore, traditional model-based defenses against MIA are infeasible. Following (Olatunji, Nejdl, and Khosla 2021), we also assume the attacker knows the target data distribution and can access a part of in-distribution data $G_{s}$ (shadow data). To conduct MIA in FGC scenarios, the attacker first utilizes $G_{s}$ to train a shadow GNN $\mathcal{S}$. Then, the posterior probabilities of nodes from $\mathcal{S}$ are used to train an attack model $\mathcal{A}$ to distinguish between non-member and member nodes. Meanwhile, the attacker utilizes $G^{\prime}$ to train a target GNN $\mathcal{T}$. During the attack phase, given a target node $v\in V$ and its $L$-hop neighbors, the posterior probability of $v$ will be obtained by feeding $v$ into $\mathcal{T}$. Then, $\mathcal{A}$ can infer the membership of $v$ based on its posterior probability. Note that the $L$-hop neighbors of $v$ known to the attacker may not be the exact neighbors, i.e., the attacker would not know the exact local structure, which is also a realistic setting.
@@ 3.1 General Framework of FGC
As illustrated in (Jin et al. 2022b), it’s hard to optimize $A^{\prime}$, $X^{\prime}$ and $Y^{\prime}$ jointly. Therefore, we fix the $Y^{\prime}$ as the same distribution as the original $Y$ during federated training. To obtain the distribution $P(Y)$, the client $j$ should upload the number of each class $n_{j}(c)$ in its own $Y_{j}$ before federated training. For updating $A^{\prime}$ and $X^{\prime}$, we model their correlations during optimizing by assuming the graph structure $A^{\prime}$ is induced from node features $X^{\prime}$: $A_{i,j}^{\prime}=\sigma(\text{MLP}_{\Phi}(x_{i}^{\prime},x_{j}^{\prime}))$, which largely reduces computation cost.
One-step communication. The challenges in FGL lie in broken local structures and distributional heterogeneity of subgraphs (Yao et al. 2023; Baek et al. 2023). In contrast, the FGC aims to learn a smaller graph from distributed subgraphs, the challenges are further compounded by the required utility-preserving. Motivated by (Yao et al. 2023), we also introduce a one-step communication between server and clients to obtain missing neighbor aggregations. This step is conducted once before federated training, and the final aggregated node embeddings are used for FGC. Note that at most 2-hop neighbors are enough in cross-silo FL (Yao et al. 2023). Thus, this process introduces little communication cost while improving performance.
@@ 3.2 Local Graph Transformation for Preserving Membership Privacy
Based on the above framework, we can condense multiple large subgraphs into a small graph without exposing local data. However, our empirical study shows that the condensed graph still exposes membership privacy. We perform FGC under the MIA introduced in the preliminary to record the utility and attack performance of the condensed graph. From Fig. 3, we can see that the performance of MIA increases along with the utility, i.e., the condensed graph will gradually leak the membership privacy of original graphs, which intuitively makes sense since it’s trained to match the local graphs. In this regard, new mechanisms are expected to condense graphs containing less membership privacy.
To achieve this, we propose to transfer the origin graph into another one, and then perform GC based on the new graph. The new graph should be equipped with two properties: (1) maintaining the utility of the original graph and (2) containing less membership privacy of the original graph. Inspired by recent work (Yang et al. 2023; Dai et al. 2023) that utilizes information bottleneck (IB) principles to extract minimal sufficient features from origin data, we propose to transform original local subgraphs with IB principles. The overall workflow is depicted in Fig. 4. Concretely, for a node in one client, we use $x\in\mathbb{R}^{d\times 1}$ to denote its original features. Then we aim to learn a information bottleneck encoder $f_{x}$ to extract minimal sufficient features $z$ of $x$:
@@
where $\mathcal{N}$ is the neighborhood subgraph of the node. According to the information bottleneck principles (Alemi et al. 2017; Wu et al. 2020), our objective function is:
@@
Upper bound of objective functions. It’s hard to directly optimize Eq.(5) due to the two MI terms. Following the information bottleneck principles (Alemi et al. 2017), we derive the upper bound of the two MI terms (The details can be found in the appendix). First, we introduce a variational distribution $q(y|z,\mathcal{N})$ to approximate $p(y|z,\mathcal{N})$, and we derive the upper bound of $-\text{I}(z,\mathcal{N};y)$:
@@
It should be noted that the typical graph information bottleneck (Wu et al. 2020) also performs graph structure transformation. Considering the joint optimization of features and structures brings burdensome computation to the clients, we only perform node feature transformation and keep the original structure unchanged.
@@
Federated training with transformed graphs. For each client $j$, we can obtain a new subgraph $G^{t}_{j}=(A_{j},Z_{j},Y_{j})$ after local graph transformation. The $Z_{j}$ with the best validation performance is stored locally. Then, the client utilizes $G^{t}_{j}$ to perform FGC using the general framework in Section 3.1. However, there exist two major flaws: (1) lack of enough labeled samples to obtain high-quality $Z_{j}$. Since each client only owns a subgraph of the entire graph, the labeled nodes in each client are scarce, making it hard to apply IB principles (Dai et al. 2023). (2) the obtained $Z_{j}$ of different clients are not in the same feature space. To better protect privacy, the graph transformation is conducted locally, i.e., there is no parameter sharing across clients, leading to a heterogeneous feature space and undermining the subsequent FGC. To address the first flaw, we propose a federated self-train strategy to expand labeled node sets. Before local graph transformation, all clients first collaboratively train a GNN model $\phi_{st}$ based on Fedavg (McMahan et al. 2017). Then, each client utilizes the global gnn model to label local unlabeled nodes. All the labeled nodes are utilized to train the graph transformation module. To address the second flaw, we propose to initialize the same $\phi$ among clients with pre-trained $\phi_{st}$ and fix it during federated training. Each client only needs to locally update $W^{(1)}$ and $W^{(2)}$. In this way, $z$ in different clients are forced to be unified into the same space.
@@ 3.3 Privacy and Utility Analysis
In our case, the attack aims to infer $\mathcal{G}_{i}$ after observing $\mathcal{G}^{\prime}$. Therefore, we should minimize $\Delta C=c^{*}_{\mathcal{G}_{i}}-\mathbb{E}_{P_{\mathcal{G}^{\prime}}}\left[c^{*}_{\mathcal{G}^{\prime}}\right]$.
@@
Lemma 1 demonstrates $\Delta C$ is upperbounded by a function of $\text{I}(\mathcal{G}_{i};\mathcal{G}^{\prime})$ under any bounded $C$.
@@
That is, minimizing $\text{I}(A_{i},X_{i};Z_{i})$ could defend the MIA theoretically based on the inference cost gain definition. As for utility, since $\mathcal{G}^{\prime}$ stems from $\mathcal{G}_{i}^{t}$, we can directly maximum the utility of $\mathcal{G}_{i}^{t}$, i.e., maximize $\text{I}(A_{i},Z_{i};Y_{i})$. Noting that this maximization will increase the lower-bound of $\text{I}(A_{i},X_{i};Z_{i})$ in light of $\text{I}(\mathcal{G}_{i}^{t};Y_{i})\leq\text{I}(\mathcal{G}_{i}^{t};\mathcal{G}_{i})$, thus the objective in Eq.(5) can be seen as a trade-off between privacy- and utility-preserving of the condensed graph, which is controlled by $\gamma$. We will present this effect in the experiments. The proofs of these analyses can be found in the Appendix.

### v4

@@ Abstract
Graph condensation (GC), which reduces the size of a large-scale graph by synthesizing a small-scale condensed graph as its substitution, has benefited various graph learning tasks. However, existing GC methods rely on centralized data storage, which is unfeasible for real-world decentralized data distribution, and overlook data holders’ privacy-preserving requirements. To bridge this gap, we propose and study the novel problem of federated graph condensation (FGC) for graph neural networks (GNNs). Specifically, we first propose a general framework for FGC, where we decouple the typical gradient matching process for GC into client-side gradient calculation and server-side gradient matching, integrating knowledge from multiple clients’ subgraphs into one smaller condensed graph. Nevertheless, our empirical studies show that under the federated setting, the condensed graph will consistently leak data membership privacy, i.e., the condensed graph during federated training can be utilized to steal training data under the membership inference attack (MIA). To tackle this issue, we innovatively incorporate information bottleneck principles into the FGC, which only needs to extract partial node features in one local pre-training step and utilize the features during federated training. Theoretical and experimental analyses demonstrate that our framework consistently protects membership privacy during training. Meanwhile, it can achieve comparable and even superior performance against existing centralized GC and federated graph learning (FGL) methods.
@@ 1 Introduction
Graph data, such as social networks, transportation networks, and molecular interactions, is ubiquitous in the real world. Graph neural networks (GNNs) have been prevalent for modeling these graphs and achieved remarkable success in recent years (Hamilton, Ying, and Leskovec 2017; Kipf and Welling 2017; Velickovic et al. 2018). However, GNNs have struggled with real-world large-scale graphs due to memory and computational overheads, which are further exacerbated when the models need to be trained multiple times, e.g., neural architecture search (NAS) (Zhang et al. 2022) and continual learning (Li and Hoiem 2016). A natural solution from the data-centric view is graph condensation (GC) (Jin et al. 2022b, a; Liu et al. 2022; Zheng et al. 2023), as shown in Figure 1(a), which reduces large-scale graphs into small ones but keeps the utility of original graphs, facilitating the storage of graph data and computations of downstream tasks. As a typical technique of GC, gradient matching (Jin et al. 2022b, a) forces the gradients of GNN models trained on the original graph and condensed graph to be the same, maintaining a comparable model performance, but largely reducing the graph size.
@@
Despite significant progress in GC, existing works all hold a basic assumption that the graph data is centrally stored. In reality, the whole graph is divided into multiple subgraphs held by different parties. For example, in healthcare systems, a hospital only possesses a subset of patients (nodes), their information (attributes), and interactions (links). Due to privacy concerns, data holders are unwilling to share their data and many strict regulations such as GDPR (Voigt and Von dem Bussche 2017) also forbid collecting data arbitrarily, which makes centralized GC unfeasible. Federated graph learning (FGL) (Yao et al. 2023; Wan et al. 2022; Zhang et al. 2021b) has emerged in recent years to endow collaboratively training a global model without exposing local data, as is shown in Figure 1(b). Typically, they are dedicated to training a global model by exchanging encrypted information for recovering missing structures (e.g., cross-client edges and neighbors). However, they still suffer from burdensome computations of traditional GNNs for large-scale graph training, especially in the cross-silo scenario (Gao, Yao, and Yang 2022). For example, if the NAS for FGL is conducted, it inevitably performs federated training multiple times, which exacerbates both computations and communications. Therefore, an important yet unexplored area lies in considering GC in the federated setting.
@@
To bridge this gap, we investigate the novel problem of federated graph condensations (FGC), as shown in Figure 1 (b), which aims to collaboratively learn a small graph that contains knowledge from different data sources. Subsequently, the condensed graph can be accessed by all data holders for downstream applications, such as node classifications. This task is non-trivial due to two significant challenges. (1) How to preserve the utility of the condensed graph that is learned from distributed subgraphs? Unlike centralized GC the graph data is integral, the entire graph is distributed as subgraphs across data holders in FGC. As a result, some critical cross-client information is missing and these subgraphs also suffer from severe heterogeneity, which inevitably degrades the utility of the condensed graph. In this regard, a unified framework for utility-preserving FGC is desired. (2) How to preserve the local graph’s membership privacy that may be leaked by the condensed graph? Membership privacy is widely used to measure information leakage about training data (Nasr, Shokri, and Houmansadr 2019). Unlike local storage of graph data in typical FGL, a condensed graph is released in FGC, which may be utilized to infer the membership privacy of local data using membership inference attack (MIA). Furthermore, the accessibility of the condensed graph enables attackers to train arbitrary models, as shown in Figure 2, rendering traditional model-based defenses against MIA infeasible. Thus, new methods are needed to defend against MIA in FGC.
@@
To tackle these challenges, we endeavor to condense graphs in the federated setting and propose a Federated Graph Condensation framework (FedGC). (1) To preserve data utility, we propose a general framework for FGC, the core of which is that the condensed graph is learned by matching the aggregated gradients from clients. Particularly, the gradients are aggregated in a class-aware weighted manner to tackle data heterogeneity. Furthermore, an encrypted one-step communication between the server and clients is adopted to recover the cross-client neighbors. During the evaluation phase, the model trained on the condensed graph will be fine-tuned by local data for better personalization. (2) To preserve membership privacy, we further propose a novel local graph transformation module with information bottleneck principles (Alemi et al. 2017). We first empirically reveal that the condensed graph will consistently leak membership privacy. To address this, each subgraph will be transformed locally before condensation by fixing the graph structure unchanged and extracting sufficient but minimal node features. Then the transformed graph will act as a proxy of the original graph to conduct FGC, thereby protecting the local membership privacy. Besides, to alleviate the performance decay in graph transformation caused by scarce labeled nodes and heterogeneous feature space, we propose a one-time self-training strategy to label nodes and utilize a fixed pre-trained shared model to align the feature space. Finally, we give theoretical and empirical analysis to prove the utility- and privacy-preserving properties of FedGC. The major contributions of this paper are summarized as follows:
@@
  • We conduct extensive experiments on five real-world datasets and show that FedGC outperforms centralized GC and FGL methods, especially in large-scale datasets. Meanwhile, FedGC can consistently protect membership privacy during the whole federated training process.
@@ 2 Preliminary
MIA on condensed graphs. Following (Dong, Zhao, and Lyu 2022), we assume an honest-but-curious server is the strong attacker. Since the condensed graph $G^{\prime}$ is publicly accessed, the target models can be arbitrarily selected and trained by attacks, which is different from traditional MIA on graphs (the attacker can only query target models (Olatunji, Nejdl, and Khosla 2021)), as shown in Figure 2. Therefore, traditional model-based defenses against MIA are infeasible. Following (Olatunji, Nejdl, and Khosla 2021), we also assume the attacker knows the target data distribution and can access a part of in-distribution data $G_{s}$ (shadow data). To conduct MIA in FGC scenarios, the attacker first utilizes $G_{s}$ to train a shadow GNN $\mathcal{S}$. Then, the posterior probabilities of nodes from $\mathcal{S}$ are used to train an attack model $\mathcal{A}$ to distinguish between non-member and member nodes. Meanwhile, the attacker utilizes $G^{\prime}$ to train a target GNN $\mathcal{T}$. During the attack phase, given a target node $v\in V$ and its $L$-hop neighbors, the posterior probability of $v$ will be obtained by feeding $v$ into $\mathcal{T}$. Then, $\mathcal{A}$ can infer the membership of $v$ based on its posterior probability. Note that the $L$-hop neighbors of $v$ known to the attacker may not be the exact neighbors, i.e., the attacker would not know the exact local structure, which is also a realistic setting.
@@ 3.1 General Framework of FGC
As illustrated in (Jin et al. 2022b), it’s hard to optimize $A^{\prime}$, $X^{\prime}$ and $Y^{\prime}$ jointly. Therefore, we fix the $Y^{\prime}$ as the same distribution as the original $Y$ during federated training. To obtain the distribution $P(Y)$, the client $j$ will upload the number of each class $n_{j}(c)$ in its own $Y_{j}$ before federated training. We also assume that $A^{\prime}$ is induced from $X^{\prime}$: $A_{i,j}^{\prime}=\sigma(\text{MLP}_{\Phi}(x_{i}^{\prime},x_{j}^{\prime}))$, which largely reduces computation cost.
One-step communication. The challenges in FGL lie in broken local structures and the distributional heterogeneity of subgraphs (Yao et al. 2023; Baek et al. 2023), which are further compounded by the utility-preserving property in FGC. Therefore, inspired by (Yao et al. 2023), we leverage homomorphic encryption (Brakerski, Gentry, and Vaikuntanathan 2014) to securely transfer missing neighbor aggregations among clients. This step is performed only once in advance, and the final aggregated node embeddings are used for FGC. Notably, as at most 2-hop neighbors are sufficient in cross-silo FL (Yao et al. 2023), this step introduces little communication cost while improving performance.
@@ 3.2 Local Graph Transformation for Preserving Membership Privacy
Based on the above framework, we can condense multiple large subgraphs into a small graph without exposing local data. However, our empirical study shows that the condensed graph still exposes membership privacy. We perform FGC under the MIA introduced in the preliminary to record the utility and attack performance of the condensed graph. From Fig. 3, we can see that the performance of MIA increases along with the utility, i.e., the condensed graph will gradually leak the membership privacy of original graphs, which intuitively makes sense since it’s trained to match the local graphs. Hereby, new mechanisms are desired to condense graphs containing less membership privacy.
To achieve this, we propose to transform the local subgraph into another one, and then perform FGC based on these transformed graphs. The new graph should be equipped with two properties: (1) maintaining the utility of the original graph and (2) containing less membership privacy of the original graph. Inspired by recent work (Yang et al. 2023; Dai et al. 2023) that utilizes information bottleneck (IB) principles to extract minimal sufficient features from origin data, we propose to transform original local subgraphs with IB principles. The overall workflow is depicted in Fig. 4. Concretely, for a node in one client, we use $x\in\mathbb{R}^{d\times 1}$ to denote its original features. Then we aim to learn an IB encoder $f_{x}$ to extract minimal sufficient features $z$ of $x$:
@@
where $\mathcal{N}$ is the neighborhood of the node. According to the IB principles (Alemi et al. 2017; Wu et al. 2020), our objective function is:
@@
Upper bound of the objective function. It’s hard to directly optimize Eq.(5) due to the two MI terms. Following the IB principles (Alemi et al. 2017), we derive the upper bound of the two MI terms. First, we introduce a variational distribution $q(y|z,\mathcal{N})$ to approximate $p(y|z,\mathcal{N})$, and we derive the upper bound of $-\text{I}(z,\mathcal{N};y)$:
@@
It should be noted that the typical graph information bottleneck (Wu et al. 2020) also performs graph structure transformation. Considering the joint optimization of features and structures brings burdensome computation to the clients, we only perform node feature transformation and keep the original structure unchanged, which can be further theoretically proven to defend against MIA.
@@
Federated training with transformed graphs. For each client $j$, we can obtain a new subgraph $G^{t}_{j}=(A_{j},Z_{j},Y_{j})$ after local graph transformation. The $Z_{j}$ with the best validation performance is stored locally. Then, the client utilizes $G^{t}_{j}$ to perform FGC using the general framework in Section 3.1. However, there exist two major flaws: (1) lack of enough labeled samples to obtain high-quality $Z_{j}$. Since each client only owns a subgraph of the entire graph, the labeled nodes in each client are scarce, making it hard to apply IB principles (Dai et al. 2023). (2) the obtained $Z$ of different clients are not in the same feature space. To better protect privacy, the graph transformation is conducted locally, i.e., there is no parameter sharing across clients, leading to a heterogeneous feature space and undermining the subsequent FGC. To address the first flaw, we propose a federated self-train strategy to expand labeled node sets. Before local graph transformation, all clients first collaboratively train a GNN model $\phi_{st}$ based on Fedavg (McMahan et al. 2017). Then, each client utilizes $\phi_{st}$ to label local unlabeled nodes. To address the second flaw, we propose to initialize the same $\phi$ among clients with pre-trained $\phi_{st}$ and fix it during federated training. Each client only needs to locally update $W^{(1)}$ and $W^{(2)}$. In this way, $Z$ in different clients are forced to be unified in the same space.
@@ 3.3 Privacy and Utility Analysis
In our case, the attack aims to infer $\mathcal{G}_{i}$ after observing $\mathcal{G}^{\prime}$. Therefore, we should minimize $\Delta C=c^{*}_{\mathcal{G}_{i}}-\mathbb{E}_{P_{\mathcal{G}^{\prime}}}\left[c^{*}_{\mathcal{G}^{\prime}}\right]$. Inspired by (Makhdoumi et al. 2014), we have:
@@
Lemma 1 demonstrates that $\Delta C$ is upperbounded by a function of $\text{I}(\mathcal{G}_{i};\mathcal{G}^{\prime})$ under any bounded $C$.
@@
Therefore, minimizing $\text{I}(A_{i},X_{i};Z_{i})$ could defend the MIA theoretically based on the definition of inference cost gain. As for utility, since $\mathcal{G}^{\prime}$ stems from $\mathcal{G}_{i}^{t}$, we can directly maximum the utility of $\mathcal{G}_{i}^{t}$, i.e., maximize $\text{I}(A_{i},Z_{i};Y_{i})$. Noting that this maximization will increase the lower-bound of $\text{I}(A_{i},X_{i};Z_{i})$ in light of $\text{I}(\mathcal{G}_{i}^{t};Y_{i})\leq\text{I}(\mathcal{G}_{i}^{t};\mathcal{G}_{i})$, thus the objective in Eq.(5) can be seen as a trade-off between privacy- and utility-preserving of the condensed graph, which is controlled by $\gamma$. We will present this effect in the experiments. The proofs of these analyses can be found in the Appendix.